if f(x)=x^2+x and g(x)=x+1 then how do you solve for (g o f)(x) as in g(f(x))?
Answer by user_dude2008(1862)

f(x)=x^2+x

f(g(x))=(x+1)^2+(x+1)

f(g(x))=x^2+2x+1+x+1

f(g(x))=x^2+3x+2